A research group conducted an extensive survey of 3186 wage and salaried workers on issues ranging from relationships with their bosses to household chores. The data were gathered through hour-long telephone interviews with a nationally representative sample. In response to the question, "What does success mean to you?" 1641 responded, "Personal satisfaction from doing a good job." Let p be the population proportion of all wage and salaried workers who would respond the same way to the stated question. How large a sample is needed if we wish to be 95% confident that the sample percentage of those equating success with personal satisfaction is within 1.4% of the population percentage? (Hint: Use p ≈ 0.52 as a preliminary estimate. Round your answer up to the nearest whole number.)

answer: _________ workers
Answer #1

Solution,

Given that,

\hat p = 0.52

1 - \hat p = 1 - 0.52 = 0.48

margin of error = E = 0.014

At 95% confidence level

\alpha = 1 - 95%

\alpha = 1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96   

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.014)2 * 0.52 * 0.48

= 4892.16

sample size = n = 4893 workers.